Circular No. 5698

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET TEMPEL 1 (1993c) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the Spacewatch 0.9-m telescope at Kitt Peak:

1993 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 21.19826  6^{h}33^{m}22.^{s}43  +31^{o}00′14′′.4  20.7
21.20728  6 33 21.95  +31 00 15.0  21.2
21.21597  6 33 21.49  +31 00 15.6  21.2
21.22474  6 33 21.02  +31 00 16.3  21.0
22.26829  6 32 26.75  +31 01 11.8  21.2
22.27729  6 32 26.30  +31 01 12.2  21.4
22.28610  6 32 25.82  +31 01 12.6  21.3

The object is of essentially stellar appearance and the position in close agreement with the prediction on MPC 18258.

(5145) PHOLUS J. Davies, Royal Observatory, Edinburgh; J. Spencer, Lowell Observa-

tory; M. Sykes, Steward Observatory; D. Tholen, Institute for Astronomy, University of Hawaii; and S. Green, University of Kent, report: “We obtained thermal infrared observations of (5145) Pholus from Mauna Kea. Following are the mean times of observation and the corresponding Q magnitudes. The first two observations were by Spencer and Tholen with the NASA Infrared Telescope Facility, and the third observation was by Davies with the U.K. Infrared Telescope (UKIRT): Jan. 11.66 UT, 3.87_{−0.34}^{+0.50}; 12.42, 4.47_{−0.34}^{+0.50}; 14.56, 3.86 ± 0.3. On Jan. 12, pointing was by blind offset from a nearby star due to high sky brightness from the moon. The minor planet was visible and had the expected visual magnitude as determined by simultaneous optical photometry. Use of a standard thermal model with the UKIRT data suggest a minimum diameter of 189 ± 26 km and a maximum albedo of 0.044 ± 0.013.”

COMET OHSHITA (1992a_{1}) H. Mikuz, Ljubljana, Slovenia, reports m_{1} = 14.4 on Jan. 19.20 UT,

using a 0.20-m Baker-Schmidt camera (+ CCD + V filter).

1993 January 27 Daniel W. E. Green